A horizontal forces F pushes a 9kg load up an inclined plane with an acceleration of 1.2m/s^2. what is the magnitude of the force if the plane makes an angle of 30° with the horizontal and the coefficient of kinetic friction between the load and the plane is 0.15?
1
Expert's answer
2020-07-02T17:17:18-0400

Let's apply the Newton's Second Law of Motion in projections on axis "x"- and "y":


"\sum F_x = ma_x = ma,""\sum F_y = ma_y = 0,""F_{appl} - mgsin\theta - F_{fr} = ma,""N - mgcos\theta = 0,""F_{appl} - mgsin\theta - \mu_k N = ma,""N = mgcos\theta,""F_{appl} - mgsin\theta - \mu_k mgcos\theta = ma."

From this equation we can find the the magnitude of the applied force:


"F_{appl} = ma + mg(sin\theta + \mu_k cos\theta),"

"F_{appl} = 9 \ kg \cdot 1.2 \ \dfrac{m}{s^2} + 9 \ kg \cdot 9.8 \ \dfrac{m}{s^2} \cdot (sin30^{\circ} + 0.15 \cdot cos30^{\circ}) = 66.3 \ N."

Answer:

"F_{appl} = 66.3 \ N."